# Optimal market-neutral multivariate pair trading on the cryptocurrency platform

### v3

@@ 3.3.5 Market Neutrality
$-w_{n,short}\cdot(1+tc)P_{c1,t}=w_{n,long}\cdot(1-tc)P_{c2,t}$
$\forall t\in T,\forall(c1,c2)\in C\;,\forall n\in N$ (5)
@@ 3.3.6 Optimization Equation
$\begin{cases}\text{obj}:&\sum_{i=1}^{|N|}w_{n}\cdot(EP_{n}\odot[1,-1])^{{}^{\prime}}\,,\,\forall n\in N\\ &-\lambda\sum_{i=1}^{|N|}w_{n}\cdot cov_{n}\odot\begin{bmatrix}1&-1\\ -1&1\end{bmatrix}\cdot w_{n}^{{}^{\prime}}\,,\,\forall n\in N\\ s.t.&0\leq w_{n,long}\leq 1,\;-1\leq w_{n,short}\leq 0,\;\forall n\in N\\ \\ s.t.&TW_{c}+\sum_{n=1}^{N}w_{c,long}-\sum_{n=1}^{N}w_{c,short}\leq 1\,,\,\forall c\in C\\ s.t.&-w_{n,short}\cdot(1+tc)P_{c1,t}=w_{n,long}\cdot(1-tc)P_{c2,t}\\ &\forall t\in T,\forall(c1,c2)\in C\,,\,\forall n\in N\end{cases}$ (6)
@@ 4 Simulation
To calculate the amount to invest, we call the optimisation function (Algorithm 1 Line 1) by feeding trading data and properties of trading data such as expected return and expected risk to calculate the weights. The optimization is based on Equation 6. Once the optimal weights are calculated, they are returned to the trading function for the continuation of position opening until the next call. The code implementing our methodology is available in a public repository [^1] The source code can be accessed at https://github.com/Hongshen-Yang/OTT.
@@ 5.1 Experimental Setup
We trade a bucket of fiat currency pairs referenced to ETH for several reasons: ① Cryptocurrency markets have substantial liquidity and depth; ② The volatility in the cryptocurrency market overwhelms the inherent volatility in fiat currencies, providing opportunities for arbitrage. ③ For fiat currency pairs (equivalently spreads) referenced to cryptocurrencies, ETH has the strongest cointegrated and stationary relationship comparing to BTC (c.f. Table 3, LABEL:tab:btc_corr);

### v4

@@ 3.3.5 Market Neutrality
$-w_{n,short}\cdot(1+tc)P_{c_{1},t}=w_{n,long}\cdot(1-tc)P_{c_{2},t}$
$\forall t\in T,\forall(c_{1},c_{2})\in C\;,\forall n\in N$ (5)
@@ 3.3.6 Optimization Equation
$\begin{cases}\text{obj}:&\sum_{i=1}^{|N|}w_{n}\cdot(EP_{n}\odot[1,-1])^{{}^{\prime}}\,,\,\forall n\in N\\ &-\lambda\sum_{i=1}^{|N|}w_{n}\cdot cov_{n}\odot\begin{bmatrix}1&-1\\ -1&1\end{bmatrix}\cdot w_{n}^{{}^{\prime}}\,,\,\forall n\in N\\ s.t.&0\leq w_{n,long}\leq 1,\;-1\leq w_{n,short}\leq 0,\;\forall n\in N\\ \\ s.t.&TW_{c}+\sum_{n=1}^{N}w_{c,long}-\sum_{n=1}^{N}w_{c,short}\leq 1\,,\,\forall c\in C\\ s.t.&-w_{n,short}\cdot(1+tc)P_{c_{1},t}=w_{n,long}\cdot(1-tc)P_{c_{2},t}\\ &\forall t\in T,\,\forall(c_{1},c_{2})\in C,\,\forall n\in N\end{cases}$ (6)
Where
  • $N$: Set of all pairs
  • $w_{n}$: Weight vector for the nth pair
  • $EP_{n}$: Expected profit vector for the nth pair
  • $\odot$: Hadamard product (element-wise multiplication)
  • $\lambda$: Risk aversion coefficient
  • $cov_{n}$: Covariance matrix for the nth pair
  • $w_{n,long}$: Long weight for the nth pair
  • $w_{n,short}$: Short weight for the nth pair
  • $TW_{c}$: Trading weight for currency $c$
  • $c$: Currency
  • $tc$: Transaction cost
  • $P_{c_{1},t}$: Price of currency $c_{1}$ at time $t$
  • $P_{c_{2},t}$: Price of currency $c_{2}$ at time $t$
  • $T$: Set of all time periods
@@ 4 Simulation
To calculate the amount to invest, we call the optimisation function (Algorithm 1 Line 1) by feeding trading data and properties of trading data such as expected return and expected risk to calculate the weights. The optimization is based on Equation 6. Once the optimal weights are calculated, they are returned to the trading function for the continuation of position opening until the next call. The code implementing our methodology is available in a public repository [^1] The source code can be accessed at https://github.com/Hongshen-Yang/optimal-trading-technique.
@@ 5.1 Experimental Setup
We trade a bucket of fiat currency pairs referenced to ETH for several reasons: ① Cryptocurrency markets have substantial liquidity and depth; ② The volatility in the cryptocurrency market overwhelms the inherent volatility in fiat currencies, providing opportunities for arbitrage. ③ For fiat currency pairs (equivalently spreads) referenced to cryptocurrencies, ETH has the strongest cointegrated and stationary relationship (c.f. Table 3);
@@ 6 Conclusion
Appendix A Sensitivity Analysis
To further examine the robustness of our Optimal Trading Technique (OTT)method, we expanded the analysis to include additional cryptocurrencies and a more extensive time period. While ETH is the most ideal candidate for the application of the OTT method, we also included BTC (Bitcoin) and SOL (Solana) as intermediates for pair trading based on their market capitalization. Given that BTC and ETH are the primary cryptocurrencies tradeable against multiple fiat currencies, we focused on USD, GBP, and EUR as the common denominators. Additionally, since the selected full-cycle, bull-run, and bear-run markets coincide with the COVID-19 period [^6] COVID-19, a pandemic since 2019, has had a significant global impact (Shi et al., 2020), we extended the experiment to the post-COVID period using the latest publicly available data from Kraken up to 01-Apr-2024. Major cryptocurrencies have generally experienced exponential growth, contributing to a strong Buy-and-Hold profitability.
Table 11: Correlation and Cointegration among Currency Pairs
BTC Interval 15min Interval 60min Interval 720min
Currencies Corr Coint Corr Coint Corr Coint
USD:GBP 0.9530 (83.7%) 0.9504 (87.5%) 0.9480 (55.8%)
USD:EUR 0.9687 (76.0%) 0.9672 (70.2%) 0.9659 (31.7%)
GBP:EUR 0.9806 (94.2%) 0.9792 (98.1%) 0.9776 (73.1%)
ETH Interval 15min Interval 60min Interval 720min
Currencies Corr Coint Corr Coint Corr Coint
USD:GBP 0.9773 (69.2%) 0.9745 (81.7%) 0.9717 (79.8%)
USD:EUR 0.9839 (56.7%) 0.9822 (70.2%) 0.9807 (37.5%)
GBP:EUR 0.9885 (88.5%) 0.9873 (95.2%) 0.9857 (94.2%)
SOL Interval 15min Interval 60min Interval 720min
Currencies Corr Coint Corr Coint Corr Coint
USD:GBP 0.9896 (76.9%) 0.9892 (86.5%) 0.9865 (95.2%)
USD:EUR 0.9932 (60.6%) 0.9929 (79.8%) 0.9918 (84.6%)
GBP:EUR 0.9958 (89.4%) 0.9956 (98.1%) 0.9936 (99.0%)
  • This table provides average rolling correlation calculations and cointegration test results for each currency pair between 01-Jan-2022 and 01-Jan-2024.
Table 11 demonstrates a relatively strong correlation among fiat currencies and a significant level of cointegration relationships. This period is characterized by a massive bull-run market, wherein the volatility among fiat currencies is negligible compared to the substantial appreciation of cryptocurrencies. Based on these observations, we will implement our OTT techniques on all three cryptocurrencies to assess their effectiveness across different market conditions.
Table 12: Annualized Profitability Results for Popular Cryptocurrencies
ETH BTC SOL
Interval BH OTT BH OTT BH OTT
15 min (TC 0.1%) 357.88% 22.66% 246.99% 18.24% 1856.17% 7.42%
15 min (TC 0.0%) 358.80% 44.76% 247.68% 41.92% 1860.08% 68.42%
60 min (TC 0.1%) 357.88% 19.01% 246.99% 38.78% 1856.17% 3.7%
60 min (TC 0.0%) 358.80% 36.75% 247.68% 43.48% 1860.08% 33.67%
720 min (TC 0.1%) 357.88% 28.01% 246.99% 28.21% 1856.17% -0.8%
720 min (TC 0.0%) 358.80% 15.72% 247.68% 31.04% 1860.08% 3.26%
  • * The arbitrage period is 01-Oct-2020 to 01-Apr-2024 (547 days).
  • * Participating assets are USD, GBP, and EUR.
  • * The annualized returns are stated in percentages.
  • * BH (Buy-Hold), TC (Transaction Cost), OTT (Optimal Trading Technique)
  • * risk-adjust $\lambda$ is set to 1.
  • * open_threshold and close_threshold are tuned to the most lucrative combination.
The experiment of OTT on ETH, BTC and SOL under different time intervals are presented in Table 12. Overall, the results demonstrated strong profitability not only for ETH but also for BTC. ETH and BTC, which dominate the cryptocurrency market, accounting for approximately 65% of the total market cap (Coingecko, 2024), showed robust results in pair trading. While SOL achieved profitability in most trading scenarios, it did not match the performance of BTC and ETH. This is likely due to SOL’s lower trading volume and higher volatility during the examined period. Despite the significant bull-run experienced by major cryptocurrencies, our OTT method consistently produced positive returns, further validating its effectiveness across different market conditions.
Finally, extending our experiment to include the post-COVID period provided additional validation for the robustness of our OTT method. Despite the unique market conditions during the pandemic, the method maintained its effectiveness, demonstrating its adaptability and resilience across different market environments. This adaptability is crucial for traders seeking consistent returns amidst varying market dynamics.